TASK TIME (Minutes)
1 Gather .7
2 Paste .6
3 Pound 1
4 Pull .9
5 Pack .3

Assuming a steady state, what is the cycle time and capacity for a batch of 10?

cycle time and capacity for batch of 100?
Answer #1

The cycle time and capacity for a batch of 10:

Cycle time = 0.7+0.6+1.0+0.9+0.3 = 3.5 minutes per unit

Total time for a batch of 10 = 3.5 * 10 = 35 minutes

Capacity = 60 minutes / 35 minutes = 1.714286 batches per hour of “Batch of 10”

The cycle time and capacity for a batch of 100:

Cycle time = 0.7+0.6+1.0+0.9+0.3 = 3.5 minutes per unit

Total time for a batch of 100 = 3.5 * 100 = 350 minutes

Capacity = 60 minutes / 350 minutes = 0.171429 batch per hour of “Batch of 100”